OpenStudy (anonymous):

Perform the indicated operation. Simplify if possible. (8x^2/x-1) + (-8x/x-1) I factored and then canceled to get... 8x(x-1)/x-1 =8x I am stuck once again...not sure if this is correct. Thanks ahead of time.

OpenStudy (anonymous):

\[\frac{8x^2}{x-1}+\frac{-8x}{x-1}\]??

OpenStudy (anonymous):

\[\frac{8x^2-8x}{x-1}=\frac{8x(x-1)}{x-1}\] yeah looks like you are right, it is \(8x\)

OpenStudy (anonymous):

you are not stuck, you are done !

OpenStudy (anonymous):

Thank you so much! i wasn't positive it was correct, so i couldn't get past it, haha. Thanks again.

OpenStudy (anonymous):

yw